# Machine Learning Should Maximize Welfare, Not (Only) Accuracy

### v1

@@ Abstract
Decades of research in machine learning have given us powerful tools for making accurate predictions. But when used in social settings and on human inputs, better accuracy does not immediately translate to better social outcomes. This may not be surprising given that conventional learning frameworks are not designed to express societal preferences—let alone promote them. This position paper argues that machine learning is currently missing, and can gain much from incorporating, a proper notion of social welfare. The field of welfare economics asks: how should we allocate limited resources to self-interested agents in a way that maximizes social benefit? We argue that this perspective applies to many modern applications of machine learning in social contexts, and advocate for its adoption. Rather than disposing of prediction, we aim to leverage this forte of machine learning for promoting social welfare. We demonstrate this idea by proposing a conceptual framework that gradually transitions from accuracy maximization (with awareness to welfare) to welfare maximization (via accurate prediction). We detail applications and use-cases for which our framework can be effective, identify technical challenges and practical opportunities, and highlight future avenues worth pursuing.
@@ 1 Introduction
As the influence of machine learning on our lives continues to grow, there is high hope that this will prove to be for the better. Given the unmatched ability of machine learning to make accurate predictions, we have, in principle, much reason for optimism. Accurate predictions have potential to improve the choices we make for ourselves, from the mundane (e.g., what to buy or where to eat) to highly consequential (e.g., what to study, how to invest, or even who to date). This can explain why machine learning has become the backbone of most recommendation systems, media services, online marketplaces, and social platforms. Better prediction can also provide better support for decisions made about us, such as which medical treatment to apply, when to offer financial aid, or who to hire. Learned models are therefore making their way, slowly but surely, into more conventional social domains such as health care, education, finance, transportation, law, and even government. Given how learning is increasingly shaping how we communicate, express ourselves, and are productive—its influence and integration are only likely to grow. This requires us to think carefully about the role we envision for machine learning in society, as it is forming now, and as we would like it to be.
@@
If we accept that machine learning has the capacity to be socially beneficial, then ideally, we would like it to help in improving outcomes for everyone. But economic theory casts doubt as to the feasibility of such an aspiration. A basic economic truism is that whenever people benefit from something, that something will become a scarce resource, over which people will then contend. Our key point, which we will argue throughout, is that learning in social context is inherently, and inevitably, about how some form of limited resource is allocated; we will give various examples of such. When there is more demand than resources, the inevitable reality is that not everyone can get a piece of the cake. The important implication is that learning algorithms become the determinants of allocation, finding themselves in a position of making decisions about who gets what (and who doesn’t). For some tasks this is intentional, but more often it is not.
In some cases, predictions support downstream decisions in which resources are clearly scarce, such as in university admissions or job hiring, or in which risk must be hedged, such as loan approval or insurance. In other cases, scarcity can be more nuanced, but is nonetheless a substantial factor; for example, digital goods are technically of unlimited supply, but which recommended items are eventually purchased, and from which producers, is naturally restricted by the (bounded) choices of users. We posit that the issue of limited resources and their allocation holds much more broadly, and applies to most predictive tasks that takes place in social context—even when there are no explicit resources, or when the true objective is to simply maximize accuracy.
@@
The problem is that conventional learning frameworks are designed to make predictions, not allocate resources. Thus, as we pursue ever-improving accuracy, we are possibly missing out (or even deteriorating) the ability to make effective and beneficial allocations. We believe that many instances in which the use of machine learning has led to undesired outcomes, even if unintentional, can be explained by this shortcoming. As we argue, the main issue with limited resources is not their limitation per se, but the fact that people have interest in securing those resources for themselves. This creates competition either between the system and its users (if they have differing interest), amongst users themselves (directly or indirectly), or both. The crux of conventional learning tools is that they fail to acknowledge the role of human agency and its implications on learning outcomes.
Our goal is to promote the study of learning algorithms that are aware of the economic context in which they operate, and are explicitly designed to promote favorable social outcomes—by accounting for resource scarcity and human agency, and how they relate. Towards this, our main thesis is that learning frameworks should be designed to innately support a notion of social welfare. In economics, welfare quantifies the overall benefit to agents in an economic system, and questions regarding the estimation, maximization, and distribution of welfare are central to the analysis of any such system. Drawing on ideas from the field of welfare economics, we propose to adopt its main principles—namely efficiency and equity—and adapt them to the purposes and needs of current machine learning practice. We discuss the potential merits of this approach, the technical and conceptual challenges it presents, the necessary steps to ensure its successful adoption, and the limitations that will likely still remain.
1.1 Setting and scope
For concreteness, we focus on the fundamental learning task of maximizing predictive accuracy by training a model on a sample set of labeled data. However, all the ideas described here are easily applicable to other learning paradigms, including today’s generative models. We follow the conventional supervised learning paradigm and assume data is in the form of feature-label pairs $(x,y)$ sampled iid from some unknown distribution ${D}$. The data we consider represents humans; we will typically think of inputs $x$ as describing individuals (e.g., loan application, resume, movie viewing history, past consumer choices), labels $y$ as describing outcomes of interest (e.g., loan return or default, job qualification, movie rating, satisfaction), and ${D}$ as representing the population.
Although our goal is to discuss how learning can promote social good, we will insist that accuracy remains an integral part of the learning objective. This is both because accuracy has value in and of itself, and because predictive methods are, and likely will continue to be, a primary tool used in practice. Thus, we seek to retool the conventional supervised learning framework towards the goal of welfare maximization through the use of accurate predictions. This will allow us to capitalize and build on existing tools, knowledge, and practice. Because we consider limited resources and their (sometimes implicit) allocation, our discussion must extend to consider policies operating on the basis of predictions, in the style of prediction policies of Kleinberg et al. (2015).
1.2 Limitations of current practice
@@
A school receives funding for providing students with targeted aid intended to help them pass an exit exam that is required for graduation. To decide who will be selected for the program, the school obtains past data on students that participated in earlier programs and trains a model to predict the likelihood of the aid being helpful. It then uses the model to determine which new students to admit.
@@
On its face, this approach seems reasonable—but of course things are not so simple: the school may wish to maximize the impact of aid, but doing so by maximizing accuracy falls short on some important aspects. Consider the following: [^1] Other issues are apparent in the example, e.g., distribution shift or causal considerations—but these are orthogonal to our concerns.
@@
  • Although the learning objective gives equal importance to all students, this does not imply that all predictions will be equally accurate. Since borderline cases are tougher to decide, decisions could improve by giving them precedence.
@@
  • Students that are selected may not necessarily choose or agree to participate. The program might find itself with very few students and unused resources.
@@
  • Contrarily, students that are interested in receiving aid are likely to compete for acceptance, given that space is limited. This can bias selection towards students that were able to ‘appear’ more qualified (or that cheated).
  • The above approach selects students on the basis of whether aid will be effective. It does not ask how beneficial it is for each of the students to pass the exam.
@@
  • Maximizing accuracy is used here as a proxy for maximizing the number of successes. This makes no consideration of who receives aid, nor of how overall benefits from aid are distributed across the body of students.
@@
The example shows that while accuracy goes some way towards welfare, there is still an inherent tension between them. Luckily, this tension has structure, which reveals several ‘blind spots’ of current machine learning practice. These include: (i) limited resources, (ii) system decisions, (ii) human agency, and (iv) welfare considerations. Our framework intends to capture this structure and exploit it for promoting welfare through predictive machine learning.
1.3 What has been done, and what is still missing
The issue of aligning learning with societal preferences is of course not new, nor is the introduction of social welfare (Shirali et al., 2024; Perdomo, 2024). We believe however that the current literature requires a unifying formal framework for this notion. Several subfields of machine learning have made advances on some fronts, but remain to lack in others. A prime example is the literature on fairness in machine learning (Dwork et al., 2012), which clearly aims to advance some form of social good, and for this has received much attention. One concern that has been voiced is that fairness focuses on making things equal, but is unable to express the benefit that results from such efforts (Heidari et al., 2018; Hu and Chen, 2020). From our perspective, fairness takes one step towards acknowledging that something is of limited supply (e.g., positive predictions), but this is not made explicit. More importantly, fairness constraints do not account for agency: they do not ask what users want, nor how much they stand to lose or gain. Another issue is that such constraints do not consider user actions; for example, predictions for school admissions may satisfy demographic parity, but if members of the minority group do not choose to apply initially, then this guarantee is broken (Horowitz et al., 2024).
One line of research that is close to ours is that of strategic learning. This includes topics like strategic classification (Brückner et al., 2012; Hardt et al., 2016) and performative prediction (Perdomo et al., 2020) which have drawn much recent interest. Strategic learning acknowledges and explicitly models user agency, and in some cases, scarce resources. And while some works discuss social aspects of learning (e.g., Milli et al., 2019), the focus remains primarily on maximizing accuracy (and to some degree, notions of stability at equilibrium). The idea of welfare is not a driving consideration, nor is it an integral part of these frameworks.
We believe there is need for an umbrella framework that ties these loose ends and incorporates them as special cases. Such a framework will also be useful for pointing out what is missing in current solutions, and what broader questions still needs addressing. Welfare in economics is rarely something that is forced on a system; rather, it should emerge as the solution to the question—‘how can we appropriately align incentives’? Another goal is therefore to provide social planners and policymakers a handle for steering learning toward preferable social outcomes, as they deem appropriate.
1.4 Paving the path to welfare-aware machine learning
@@
Machine learning offers powerful tools for training accurate predictive models, but the above limitations suggest that this might not suffice if we wish to solve more stressing societal problems. Not by chance, social welfare offers a complementary perspective that enables such considerations. We believe that current machine learning is missing, and could much benefit from, incorporating a notion of welfare.
Position: In social contexts, machine learning should enable us to improve our collective well-being. This requires a learning framework that is capable of expressing, and provides means to optimize, the welfare of its users. To promote this goal, we should exploit the power of learning to obtain accurate predictions—not shy away from it. But accuracy cannot be the only consideration; to support constructive social outcomes, learning must: (i) explicitly consider resources, scarcity, and allocations; (ii) account for human agency by modeling what users want, know, and do; and (iii) adhere to well-specified social welfare objectives designated by a social planner.
Throughout the paper we will make these ideas clearer, more precise, better grounded, and placed in context.
@@
Our main thesis is that machine learning can benefit from the ideas and tools that welfare economics has to offer; Sec. 2 discusses its main principles, and highlights the potential for synergy with machine learning. In Sec. 3 we present our proposed framework for welfare-maximizing machine learning and details its three orders. Sec. 4 presents alternative views, and Sec. 6 gives concluding remarks and a look to the future.
2 Welfare economics and what it can teach us
Welfare economics is the subfield of economics that is concerned with the characterization, evaluation, and maximization of social welfare in economies and societies. The main principles of welfare economics date back to Adam Smith (1759, 1776), but its formal foundations were laid out only a century later by notable neoclassic economists such as Edgeworth (1881), Marshall (1890), Pareto (1906), and Pigou (1920). Modern welfare economists include influential figures such as Hicks (1939), Arrow (1951), Sen (1970), and Stiglitz (2012)—all of which received the Nobel Prize for their contributions to this field. Given its rich history, we believe that machine learning has much to gain from adopting ideas and perspectives from this well-established discipline.
2.1 Welfare economics: crash course
@@
Efficiency requires an ability to optimize economic outcomes to a point where social benefit is maximal (in the Pareto sense). Markets are a classic example of how the actions of many self-interested agents can combine to produce efficient outcomes (Arrow and Debreu, 1954). However, there are typically many states that are maximally beneficial, but that differ in how benefits are distributed across individuals; in regard to this, markets are mostly silent. As such, equity makes a statement about the relative preference ordering over all possible states, and considers means for steering towards preferable ones. A canonical example is income distribution: all governments likely seek higher overall income (efficiency), but may disagree about whether high inequality should be permitted or suppressed (equity).
@@
In welfare economics, the primary tool for defining and promoting equity is the social welfare function, which ranks or scores all possible economic states. Our focus will be on the common choice of cardinal (i.e., real-valued) welfare functions that take the form of an expectation over weighted individual utilities: [^2] The literature on social welfare functions is of course rich and diverse. Here we present a basic and simplified formulation which suits our purposes. Curious readers are referred to (Adler, 2019) for a more thorough discussion of possible alternatives.
@@
where ${\pi}$ is a policy (e.g., rules deciding which resources are allocated to whom) and $u(x,y;{\pi})$ is the utility of user $x$ under policy $\pi$, given $y$. Notably, $w(x)$ is a weight function that the planner chooses: this defines the desired direction for overall improvement (i.e., efficiency) by balancing the importances of individual outcomes (i.e., equity). [^3] More generally, weights $w(x)$ can even depend on the utility $u$ of $x$. This allows to express welfare using broader functional forms (e.g., $\min$ instead of sum) or relative measures of inequality (e.g., Gini or Theil). For our purposes, the simpler form $w(x)$ suffices. When the policy ${\pi}={\pi}_{h}$ is guided by a predictive model $h$, we will write ${\mathtt{welfare}}(h)$ to mean ${\mathtt{welfare}}({\pi}_{h})$.
@@
Equity is inherently a subjective notion that requires making value judgements. Social welfare functions make it possible to formally express these by setting appropriate weights. In welfare economics, weights are designated by a social planner—either a real or fictitious entity that represents societal preferences. A social planner can set weights to aid low income individuals; implement affirmative action towards some social group; or ensure that all individuals obtain some minimal level of utility. Concrete examples include using weights to prioritize certain subgroups (Björkegren et al., 2022) or balance different objectives (Rolf et al., 2020). The simplest weighing scheme is of course using uniform weights, i.e., $w(x)=1$ for any $x$. But note even this makes a statement, which is that individuals should be weighted by their utility; this is known as ‘utilitarian welfare’. Hence, from the perspective of welfare economics, any objective that optimizes a (non-weighted) average—such as accuracy in machine learning—is in effect making a statement about how value should be distributed.

### v2

# but Not by (Only) Maximizing Accuracy

@@ Abstract
Decades of research in machine learning have given us powerful tools for making accurate predictions. This has made such tools appealing for use in social settings and on human inputs. Yet despite a lack of justification for why the generic approach of accuracy maximization can or should improve our collective well-being—and mounting evidence of likely adverse outcomes—it remains the widespread default. This position paper asserts that for machine learning to become socially beneficial, it must be embedded within a broader economic framework that explicitly aims to maximize social welfare. The field of welfare economics asks: how should we allocate limited resources among self-interested agents to maximize overall benefits? We contend that this perspective applies to many contemporary applications of machine learning in social contexts, and advocate for its adoption. Rather than disposing of prediction, we propose to leverage this forte of machine learning towards welfare maximization. We demonstrate this idea by portraying a conceptual framework that gradually transitions from accuracy maximization (with awareness to welfare) to welfare maximization (via accurate prediction). We detail applications and use-cases for which this framework can be effective, identify technical challenges and practical opportunities, and highlight future avenues worth pursuing.
Keywords Welfare-maximizing machine learning  $\cdot$ ML in social contexts
@@ 1 Introduction
As the influence of machine learning on our lives continues to grow, there is high hope that this will prove to be for the better. Given the unmatched ability of machine learning to make accurate predictions, we have, in principle, much reason for optimism. Accurate predictions have potential to improve the choices we make for ourselves, from mundane (e.g., what to buy or where to eat) to highly consequential (e.g., what to study, how to invest, or even who to date). This can explain why machine learning has become the backbone of most recommendation systems, media services, online marketplaces, and social platforms. Better prediction can also provide better support for decisions made about us, such as which medical treatment to apply, when to offer financial aid, or who to hire. Learned models are therefore making their way, slowly but surely, into more conventional social domains such as health care, education, finance, transportation, law, and even government. Given how learning is increasingly shaping the ways we communicate, express ourselves, and are productive—its influence and integration are only likely to grow. This requires us to think carefully about the role we envision for machine learning in society, as it is forming now, and as we would like it to be.
@@
If we accept that machine learning has the capacity to be socially beneficial, then ideally, we would like for it to help improve outcomes for everyone. But economic theory casts doubt as to the feasibility of such an aspiration. A basic economic truism is that whenever people benefit from something, that something will become a scarce resource, over which people will then contend. Our key point, which we will argue throughout, is that learning in social context is inherently, and inevitably, about how some form of limited resource is allocated; we will give various examples. And when demand exceeds available resources, the inevitable reality is that not everyone can get a piece of the cake.
@@
As learning algorithms increasingly influence who gets what and who doesn’t—whether by design or by consequence—it becomes imperative to understand the implications of learning under scarcity. The challenge is that conventional learning frameworks are innately designed to make predictions, not allocate resources. Thus, as we pursue ever-improving accuracy, we are possibly missing out on (or even deteriorating) the ability to make socially effective and beneficial allocations. The main challenge of allocation is not the limitation of resources per se, but the fact that people have interest in securing those resources for themselves. This creates competition either between the system and its users (when their interest differ), amongst users themselves (directly or indirectly), or with other stakeholders. Thus, a second crux of conventional learning is that it fails to acknowledge the role of human agency and decision-making through its implications on learning outcomes. We believe such economic grounds underlie (to some degree) many instances of machine learning leading to undesired (if unintentional) outcomes, including: perpetuated biases (Zou and Schiebinger, 2018) and group exclusion (De-Arteaga et al., 2019); Goodhart’s law (Manheim and Garrabrant, 2018) and gaming behavior (Hardt et al., 2016); content polarization (Bakshy et al., 2015), extremification (Ribeiro et al., 2020), and echo chambers (Cinelli et al., 2021); popularity bias (Abdollahpouri et al., 2017), exposure bias (Salganik et al., 2006), algorithmic bias (Kleinberg et al., 2018), and the Matthew effect (Perc, 2014) as some examples.
Rather than hoping for social outcomes to miraculously improve as a result of accuracy maximization, this position paper advocates for the study, design, and deployment of learning algorithms that directly account for the economic context in which they operate. Economic awareness does not require reinventing the entire learning pipeline; rather, we assert this can, and should, be achieved by augmenting existing predictive methodologies to support key economic notions, foremostly resource scarcity and human agency. But “better social outcomes” is an illusive concept that often evades proper definition; or when it is defined, remains constrained to a particular narrow formalism. To maintain generality and provide flexibility, we propose to build on the notion of social welfare as a core guiding principle. In economics, social welfare quantifies the overall benefit to agents in an economic system, and questions regarding the estimation, maximization, and distribution of welfare are central to the analysis of any such system. Adopting a welfare economics perspective offers many benefits: an established conceptual framework for reasoning about social outcomes, formal structure that enables effective integration with machine learning, and a practical toolkit for guiding system designers and policymakers.
Although the concept of welfare has been considered to some extent in machine learning (e.g., Rambachan et al., 2020), it has so far been limited mostly to works at the intersection of computational game theory, and in a restricted manner, to fairness in machine learning. As more and more ML sub-communities are finding themselves in a position where their work begins to have significant social impact, this paper calls for a more comprehensive interface with welfare: in theory and in practice, throughout the entire learning pipeline, and by any and all in the general learning community. Towards this, we suggest a conceptual framework consisting of three "orders" of problem formulations, targeting increasingly complex economic constructs and providing different levels of abstraction. We intend for our framework to serve as a roadmap for researchers interested in integrating welfare considerations into learning tasks of all types; to enable a bidirectional flow of ideas and methodologies between learning and economics; and to permit effective communication across the different learning communities. We discuss the potential merits of this framework, the technical and conceptual challenges it presents, the necessary steps to ensure its successful adoption, and the limitations that will likely still remain.
1.1 On the limitations of accuracy: an illustrative example
@@
A school receives funding for providing students with targeted aid intended to help them pass an exit exam that is required for graduation. To decide who will be selected for the program, the school obtains past data on students that participated in earlier programs and trains a classifier to predict the likelihood of aid being helpful. It then uses the model to determine which new students to admit.
@@
On its face, this approach seems reasonable—but of course things are not so simple: the school may wish to maximize the impact of aid, but doing so by maximizing accuracy as a proxy falls short on several important aspects. Consider the following: [^1] The example exhibits other issues, e.g., data shift and causality—which are important, but orthogonal to our goal.
@@
  • Although the learning objective gives equal importance to all students, this does not imply that all predictions will be equally accurate. Cases may differ in their influence on final decisions, their difficulty of resolution, their impact on individuals, or their implications for the entire program.
@@
  • Even if a student is selected, he or she may not necessarily choose or agree to participate. Without factoring participation, the program might find itself left with fewer or less qualified students.
@@
  • Contrarily, students interested in receiving aid are likely to compete for acceptance, given that space is limited. This can bias selection towards students trying to ‘appear’ more qualified (or cheating).
@@
  • Maximizing accuracy is used here as a proxy for maximizing the number of successes. This makes no consideration of who receives aid, nor of the distribution of benefits from aid across all students.
  • The above approach selects students on the basis of whether aid will be effective. It does not ask how beneficial it is for each of the students themselves to be admitted or to pass the final exam.
@@
The example shows that while accuracy goes some way towards welfare, there is still an inherent tension between them. Luckily, this tension has structure, which reveals several ‘blind spots’ of current machine learning practice. These include: (i) limited resources, (ii) human agency, and (iii) policy considerations. Not by chance, these are precisely the building blocks of welfare economics (see Sec. 2).
1.2 Machine learning for social welfare
@@
Machine learning offers powerful tools for training accurate predictive models, but the above limitations suggest that this might not suffice once learning encounters human reality. Social welfare offers a complementary perspective that enables such considerations. We believe that current machine learning is missing, and could much benefit from, incorporating a notion of welfare.
Position: Machine learning has the potential to improve our collective well-being, and should be used towards this purpose. But we cannot assume social good will spontaneously materialize just by improving accuracy. Nonetheless, we can, and should, rely on the power of existing predictive methodologies. Through a revised systemic perspective and with appropriate adaptation, conventional learning frameworks should be retooled with a capacity to express and optimize social welfare. This requires providing explicit support for key economics notions: resource scarcity and allocation, human agency and interactions, decision-making, and the role of policy.
We believe it is high time to both confront the challenges and seize the opportunities that learning in social context presents in a disciplined and engaging manner. Welfare considerations now apply to the machine learning community at large, and cannot remain an exclusive niche to work on. Already we are seeing ideas from economics leaking into current machine learning practice (e.g., LLM alignment by preference elicitation); rather than wait for these to surface as grassroots phenomena lacking guidance or a unified conceptualization, we should remain one step ahead and preemptively lay out the necessary groundwork in anticipation of future challenges.
We envision a welfare-centric learning paradigm that is accessible to all, simple to communicate, and appealing to work on—as are most successful learning frameworks. Welfare should be promoted by exploiting the power of learning to obtain accurate predictions, not shying away from it. This can be accomplished by merging the foundational concepts of welfare economics with the established theory and practical know-how of modern machine learning.
@@
Our main thesis is that machine learning can benefit from the ideas and tools that welfare economics has to offer; Sec. 2 discusses its main principles, and Sec. 3 highlights the potential for synergy with machine learning. In Sec. 4 we present our proposed framework for welfare-maximizing machine learning and details its three orders. Sec. 6 concludes with a look to the future.
2 Welfare economics: crash course
Welfare economics is a field concerned with the characterization, evaluation, and maximization of social welfare in economies and societies. The main principles of welfare economics date back to Adam Smith Adam Smith (1759, 1776), but its formal foundations were laid out only a century later by notable neoclassic economists such as Edgeworth (Edgeworth, 1881), Marshall (Marshall, 1890), Pareto (Pareto, 1906), and Pigou (Pigou, 1920). Modern welfare economists include influential figures such as Hicks (Hicks, 1939), Arrow (Arrow, 1951), Sen Sen (1970), and Stiglitz (Stiglitz, 2012)—all of which received the Nobel Prize for their contributions to this field. Given its rich history, we believe there is much to gain from adopting ideas and perspectives from this well-established discipline.
@@ 2 Welfare economics: crash course
Efficiency requires an ability to optimize economic outcomes to a point where social benefit is maximal (in the Pareto sense). Markets are a classic example of how the actions of many self-interested agents can combine to produce efficient outcomes (Arrow and Debreu, 1954). However, there are typically many states that are maximally beneficial, but that differ in how benefits are distributed across individuals; in regard to this, markets are mostly silent. As such, equity makes a statement about the relative preference ordering over all possible states, and considers means for steering towards preferable ones. A canonical example is income distribution: all governments likely seek higher overall income (efficiency), but may disagree about whether high income inequality should be permitted or suppressed (equity).
@@
In welfare economics, the primary tool for defining and promoting equity is the social welfare function, which ranks or scores all possible economic states. Welfare functions come in various shapes and colors; for concreteness, we focus on the common choice of cardinal (i.e., real-valued) welfare functions that take the form of an expectation over weighted individual utilities:
@@
where ${\pi}$ is a policy (e.g., rules deciding which resources are allocated to whom) and $u(x,y;{\pi})$ is the utility of user $x$ under policy $\pi$, given $y$. Notably, $w(x)$ is a weight function that the planner chooses: this defines the desired direction for overall improvement (i.e., efficiency) by balancing the importances of individual outcomes (i.e., equity). [^2] More generally, weights $w(x)$ can even depend on the utility $u$ of $x$. This allows to express welfare using broader functional forms (e.g., $\min$ instead of sum) or relative measures of inequality; see Appendix A. When the policy ${\pi}={\pi}_{h}$ is guided by a predictive model $h$, we will write ${\mathtt{welfare}}(h)$ to mean ${\mathtt{welfare}}({\pi}_{h})$. See Appendix A for further examples.
@@
Equity is inherently a subjective notion that requires making value judgements. Social welfare functions make it possible to formally express these by setting appropriate weights. In welfare economics, weights are designated by a social planner—either a real or fictitious entity that represents societal preferences. A social planner can set weights to aid low income individuals; implement affirmative action towards some social group; or ensure that all individuals obtain some minimal level of utility. Concrete examples include using weights to prioritize certain subgroups (Björkegren et al., 2022) or balance different objectives (Rolf et al., 2020). The simplest weighing scheme is of course using uniform weights, i.e., $w(x)=1$ for any $x$. But note even this makes a statement, which is that individuals should be weighted by their utility; this is known as ‘utilitarian welfare’. From the perspective of welfare economics, any objective over individuals makes some statement about how value should be distributed.